True or false ?
1. If X1, X2, ..., XE is a random sample, then
the sample mean is an unbiased estimator of
the population mean.

2. A maximum likelihood estimator is always
unbiased.
1
Expert's answer
2020-09-10T17:36:36-0400

Answer 1)


Given that X1, X2, X3,.....XN are independently selected random variables from the population with the mean "\mu" and variance "\sigma^2"


Then the expected value of the sample mean is given by,


"E[\bar X]=E[\frac {X1+X2+X3+X4.....Xn}{n}]=E[\frac {1}{n}(X1+X2+X3.....Xn)]"


"E[\bar X]=\frac {1}{n}E[X1+X2+X3+....+Xn]"


"E[\bar X]=\frac{1}{n}[E[X1]+E[X2]+E[X3]+....E[Xn]]"


But since the expected value of each of the above random samples is "\mu" we can write


"E[\bar X]=\frac {1}{n}[\mu+\mu+\mu+......\mu]=\frac {1}{n}[n*\mu]=\mu"


This implies that,


If X1, X2, ..., XE is a random sample, then the sample mean is an unbiased estimator of

the population mean = TRUE


Answer 2)


Let us check whether the sample variance "s^2" is an unbiased maximum likelihood estimator for population variance


"s^{2}=\frac{1}{N} \sum_{i=1}^{N}\left(x_{i}-\bar{x}\right)^{2}"


In statistics, we evaluate the “goodness” of the estimation by checking if the estimation is “unbiased”. By saying “unbiased”, it means the expectation of the estimator equals to the true value, e.g. if E[x] = µ then the mean estimator is unbiased. Now we will show that the equation actually holds for mean estimator.

"\begin{aligned}
\mathbb{E}[\bar{x}] &=\mathbb{E}\left[\frac{1}{N} \sum_{i=1}^{N} x_{i}\right]=\frac{1}{N} \sum_{i=1}^{N} \mathbb{E}[x] \\
&=\frac{1}{N} \cdot N \cdot \mathbb{E}[x] \\
&=\mathbb{E}[x]=\mu
\end{aligned}"

The first line makes use of the assumption that the samples are drawn i.i.d from the true distribution, thus E[xi ] is actually E[x]. From the proof above, it is shown that the mean estimator is unbiased.


Now we move to the variance estimator. At the first glance, the variance estimator "s^{2}=\frac{1}{N} \sum_{i=1}^{N}\left(x_{i}-\bar{x}\right)^{2}" should follow because mean estimator x is unbiased. However, it is not the case


"\begin{aligned}
\mathbb{E}\left[s^{2}\right] &=\mathbb{E}\left[\frac{1}{N} \sum_{i=1}^{N}\left(x_{i}-\bar{x}\right)^{2}\right] \\
&=\frac{1}{N} \mathbb{E}\left[\sum_{i=1}^{N} x_{i}^{2}-2 \sum_{i=1}^{N} x_{i} \bar{x}+\sum_{i=1}^{N} \bar{x}^{2}\right]
\end{aligned}"


"\text { We know } \sum_{i=1}^{N} x_{i}=N \cdot \bar{x} \text { and } \sum_{i=1}^{N} \bar{x}^{2}=N \cdot \bar{x}^{2} . \text { Plug these into the derivation: }"

"\begin{aligned}
\mathbb{E}\left[s^{2}\right] &=\frac{1}{N} \mathbb{E}\left[\sum_{i=1}^{N} x_{i}^{2}-2 N \cdot \bar{x}^{2}+N \cdot \bar{x}^{2}\right] \\
&=\frac{1}{N} \mathbb{E}\left[\sum_{i=1}^{N} x_{i}^{2}-N \cdot \bar{x}^{2}\right] \\
&=\frac{1}{N} \mathbb{E}\left[\sum_{i=1}^{N} x_{i}^{2}\right]-\mathbb{E}\left[\bar{x}^{2}\right] \\
&=\mathbb{E}\left[x^{2}\right]-\mathbb{E}\left[\bar{x}^{2}\right]
\end{aligned}"

"\begin{aligned}
\mathbb{E}\left[s^{2}\right] &=\left(\sigma_{x}^{2}+\mu^{2}\right)-\left(\sigma_{\bar{x}}^{2}+\mu^{2}\right) \\
&=\sigma_{x}^{2}-\sigma_{\bar{x}}^{2} \\
\sigma_{\bar{x}}^{2}=\operatorname{VAR}[\bar{x}]=& \operatorname{VAR}\left[\frac{1}{N} \sum_{i=1}^{N} x_{i}\right]=\frac{1}{N^{2}} \operatorname{VAR}\left[\sum_{i=1}^{N} x_{i}\right]
\end{aligned}"


"\operatorname{VAR}\left[\sum_{i=1}^{N} x_{i}\right]=\sum_{i=1}^{N} \operatorname{VAR}[x]=N \cdot \operatorname{VAR}[x]"


"\sigma_{\bar{x}}^{2}=\frac{1}{N} \operatorname{VAR}[x]=\frac{1}{N} \sigma_{x}^{2}"


"\mathbb{E}\left[s^{2}\right]=\frac{N-1}{N} \sigma_{x}^{2}"


"\mathbb{E}\left[s^{2}\right] \neq \sigma_{x}^{2}"


Hence we can say that the following statement,


A maximum likelihood estimator is always

unbiased = FALSE